OpenStudy (jadenwalker76):

A rectangle has a length of the cube root of 81 inches and a width of 3 to the 2 over 3 power inches. Find the area of the rectangle.

OpenStudy (jadenwalker76):

3 to the 2 over 3 power inches squared 3 to the 8 over 3 power inches squared 9 inches squared 9 to the 2 over 3 power inches squared

OpenStudy (welshfella):

area = length * width = 81 ^ (1/3) * 3 ^(2/3)

OpenStudy (welshfella):

is that correct?

OpenStudy (jadenwalker76):

yes thank you very much i first i thought it was 3^ 8/3

OpenStudy (welshfella):

well you need to make the base number = 3 81 = 3^4 so 81^(1/3) = 3^(4/3 so we have 3^4/3 * 3 ^ 2/3 now you add the powers = 3 ^ 6/3 = 3^2

OpenStudy (welshfella):

thats it

OpenStudy (jadenwalker76):

Thank u